OpenStudy (anonymous):

find an equation of the line having the given slope and containing the given point, then simplify m=5, (3,2)

OpenStudy (anonymous):

y-2 = 5(x-3) y = 5x-15+2 y = 5x-13

OpenStudy (anonymous):

y-2=5(x-3) y= 5x-15 y= 5x-15+2 y=5x -13

OpenStudy (anonymous):

need it in slope-intercept form

OpenStudy (anonymous):

It is in slope intercept form

OpenStudy (anonymous):

yes It is in slope intercept form

OpenStudy (anonymous):

5x is slope

OpenStudy (anonymous):

yeah it is. Sorry my bad. My eyes are really tired

OpenStudy (anonymous):

give us a medal

OpenStudy (anonymous):

u bet

OpenStudy (anonymous):

ty